OpenStudy (anonymous):

would i do 9P5 like i would a combination?

OpenStudy (lgbasallote):

kinda close... \[9C5 = \frac{9!}{5!(9-5)!}\] but \[9P5 = \frac{9!}{(9-5)!}\] see the difference?

OpenStudy (anonymous):

gotcha, thank you

OpenStudy (lgbasallote):

welcome